# Physics-informed machine learning for combustion: A review

### v2

@@ 4.3 Neural ODEs: surrogate models and kinetics discovery
Compared to discrete approaches, NODEs can better learn the continuous dynamics of a system, thus showing better reconstruction and extrapolation performance 291. Since NODE is proposed for learning the first-order dynamics of sequential systems, it is natural to apply it to chemical kinetic systems. Extensive studies have applied NODEs to the modeling of chemical kinetics for both building surrogate models and discovering kinetics, where many of them have informed/embedded physics to the NODEs.
@@ 4.3.1 Continuous surrogate models
As mentioned in Section 4.2, NODEs can be viewed as continuous surrogate models of chemical kinetics following the paradigm of Equation (4.14). Owoyele et al. 57 first applied NODEs to model chemical kinetics, termed as ChemNODE and shown in Figure 12. To improve the training stability, a progressive training strategy was employed, where the species were trained sequentially. For a homogeneous hydrogen-air reactor, ChemNODE showed accurate results across various operating conditions and yielded a speedup of 2.3x over the full chemical mechanism method. Kim et al. 292 used NODEs to learn stiff ODEs. The tests on the ROBER and POLLU problems demonstrated that the scaling of equations and loss functions was essential for NODEs to learn stiff ODEs.
Figure 12: The ChemNODE approach. $\bm{\psi}$ indicates the thermochemical scalars. Reprinted from Reference 57, with permission.
@@
Although the vanilla NODEs are purely data-driven, their interpretability can be enhanced by using the KANs as the network architecture. Koenig et al. 294 proposed the KAN-ODEs, combining the advantage of NODEs that do not require prior knowledge and the advantages of KANs that have increased interpretability and faster neural scaling laws compared to MLPs. The KAN-ODE was validated on learning Lotka-Volterra ODEs, modeling source terms of the Fisher-KPP PDE, and learning 1D PDEs. The scaling performance was shown superior to MLPs and the model performance could be further enhanced by sparsification and using symbolic regression. Subsequently, based on the KAN-ODEs and LeanKAN 295, Koenig et al. 283 proposed the ChemKAN for modeling and acceleration of combustion chemical kinetics. The physics was informed in two ways. First, the architecture strongly mimicked the chemical kinetics, as illustrated in Figure 13. A KAN was first used to represent the species production rates (core structure), then they were linearly combined to represent the temperature rate (thermodynamic superstructure) with another KAN to reflect the thermophysical parameter variation. This serial architecture offers the flexibility that one can only learn the species when temperature is not a variable. Second, an optional loss function for element conservation was introduced, which is similar to the PC-NODE and Phy-ChemNODE. The first test case was a biodiesel production modeling, where the thermodynamic superstructure and physical loss were "turned off". Compared with DeepONet, ChemKAN exhibited stronger scaling behaviour on test sets, less tendency toward overfitting, and stronger robustness to noise. In the second case of H2-air combustion, the ability of ChemKAN for computation acceleration was validated. Compared to MLP-based ChemNODE, ChemKAN showed similar speed-up ratios, but solved more species.
Figure 13: The structure of ChemKAN. Reprinted from Reference 283, with permission.
@@
In addition to inform physics in the form of loss functions, some studies embed the information of stoichiometry and/or rate expressions into the NODEs to enhance the physical consistency. According to Equation (3.29), the rates of species and reactions are linearly related by the stoichiometric matrix. Thus, if the stoichiometric matrix is known, one can use NODEs to directly predict $\mathbf{RoP}$ and then transform it to $\dot{\bm{\upomega}}$, or vice versa. Based on the known information, Sorourifar et al. 120 developed two versions of physics-enhanced NODEs for chemical reaction systems: r-NODE and k-NODE. The r-NODE approximates $\mathbf{RoP}$ when only the stoichiometry is known and embedded, while the k-NODE approximates the kinetic constants when the mass-action parts of $\mathbf{RoP}$ are further known and embedded. Results suggest that a ground truth model that accurately reflects the data can be learned only when the embedded information accurately reflects the underlying mechanics. Fedorov et al. 121 developed the kinetics-constrained NODEs (KC-NODEs) for kinetic modeling, where the kinetic and thermodynamic knowledge were embedded into the NN architecture, including the stoichiometric matrix, the law of mass action, and the Arrhenius dependency of the rates on temperature. The KC-NODE was successfully applied to the kinetic modeling of the hydrogenation of CO2 to methane using numerical data and to higher hydrocarbons using real experimental data. Kircher et al. 122, 123 proposed the global reaction neural network (GRNN) to learn kinetics from reactor data, where the stoichiometry and thermodynamics were explicitly embedded as the network layers, as shown in Figure 14. The GRNN can be trained directly when source data are provided and trained in a NODE way when only reactor data are available, which is a 1D plug flow reactor model in the study. It was demonstrated that the embedded knowledge enhanced the model robustness to significant noise in the data.
Figure 14: GRNN architecture with embedded stoichiometry and thermodynamics. Reprinted from Reference 122.
@@ 4.3.2 Kinetics discovery
In terms of embedding the knowledge of chemical kinetics into NNs, the chemical reaction neural network (CRNN), proposed by Ji et al. 78 for kinetics discovery, is an earlier work than References 120, 121, 122 and is also trained in a NODE-way. As shown in Figure 15, a neuron of CRNN represents a single-step reaction, where the law of mass action (Equation (3.28)) and Arrhenius law (Equation (3.32)) are embedded as the nodes and weights. For the input layer, the weights represent the reaction orders and the bias represents the rate constant in the logarithmic scale. The weights of the output layer correspond to the stoichiometric coefficients. A multi-step reaction system can be represented by stacking the neurons to formulate a CRNN with one hidden layer, where the number of hidden neurons is equal to the number of reactions. Due to the physical interpretability, once the CRNN has been trained, the kinetic parameters can be easily recovered by checking the weights and biases, then the reaction pathways can be fully revealed. The number of hidden nodes/reactions, which is a hyperparameter, can be determined by grid searching and a threshold pruning strategy is suggested to clip the NN weights below a certain threshold 78. The CRNN was first tested on three cases with increasing complexity: an elementary reaction network without temperature dependence, biodiesel production with temperature dependence, and an enzyme reaction network in which some species are present in both reactants and products. Results demonstrated that CRNN could accurately learn the reaction pathways and kinetic parameters from noisy data and even from incomplete data with missing species.
Figure 15: Schematic of the CRNN illustrated for a reaction system with four species and four reaction steps. Reprinted from Reference 78, with permission.
@@ 5 PIML for combustion reacting flows
Choi 305, 2022.1 3D+t Inverse Mass, $\mathbf{U}$, $T$, $Y_{1-3}$ No Van de Vusse Sim. MLP
@@
Sitte 306, 2022.8 2D+t Inverse Mass, $\mathbf{U}$ No Absent Sim. MLP
Liu 307, 2023.9 1D, 2D, 2D+5 Forward Mass, $\mathbf{U}$, $T$, $Y_{1-5}$ No 1-step CH4 Sim. MLP
Song 308, 2024.10 2D Both Mass, $\mathbf{U}$, $Z$, $c$ No Detailed CH4 (FPV table) Sim. MLP
Cao 309, 2024.11 2D Forward Mass, $\mathbf{U}$, $T$, $Y_{1-5}$ No 1-step CH4 Sim. MLP
@@
Peng 311, 2024.4 2D+t Inverse Mass, $\widetilde{\mathbf{U}}$, $\widetilde{T}$, $k$, $\varepsilon$ Mean Absent Sim. MLP
Liu 310, 2023.12 3D+t Inverse Mass, $\mathbf{U}$, $T$, $_{1-6}, Y\sum Y$ Direct 2-step CH4 Sim. MLP
Liu 304, 2025.9 3D+t Inverse $\mathbf{U}$ Direct Absent Sim. MLP
Taassob 312, 2023.10 2D+1 Inverse Mass, $\widetilde{\mathbf{U}}$, $\widetilde{\mathrm{PC}}_{1-2}$, $\widetilde{Z}$, $\mu_{t}$ Mean (learned) Learned Exp. MLP
@@
Yadav 313, 2024.12 (1) 1D (2)(3) 2D Inverse Mass, (1) $c$ ($\mathbf{U}$) (2) $c$ (3) $\widetilde{c}$ (1)(2) No (3) Mean (learned) (1)(2) Pre-learned (3) EBU model (1)(2) Sim. (3) Exp. MLP
Zhang 303, 2025.9 2D(+t) Inverse Mass, $\mathbf{U}$, $T$, $_{1-5}, Y\sum Y$ (1) No (2) Mean (learned) 1-step CH4 and learned (1) Sim. (2) Exp. MRF-CNN
Wang 314, 2023.7 1D+t Inverse Mass, $u$, $E$, $c$ No Simplified FRC Sim. MLP
Deng 315, 2024.7 2D+t+2 Inverse Mass, $\widetilde{\mathbf{U}}$ Mean Absent Sim. ResUNet
@@ 5.2 Laminar combustion
This subsection will present a chronological overview of studies on PINNs for laminar subsonic combustion. Choi et al. 305 used PINNs to reconstruct the 3D transient fields of a continuous stirred tank reactor (CSTR) with van de Vusse reaction. The governing equations of two reference frames were considered to enable the reconstruction of the multi-reference frame system. The fields can be accurately reconstructed from the data of velocity, temperature, and species, and some proposed strategies, such as the similarity-based sampling, have been shown to improve the model performance. Wang et al. 263 used PINNs to forward solve the $Z$-equation for a jet diffusion flame and obtained accurate results. Because the assumptions of unity Lewis number and infinite fast chemistry are adopted, the $Z$-equation is free of reaction source terms and the temperature and species can be determined solely by $Z$. Sitte et al. 306 used PINNs to reconstruct the velocity fields in puffing pool fires based on the observations of density, pressure, and temperature, following the HFM framework 247, as illustrated in Figure 16. The continuity and momentum equations are taken as the physical losses and a two-stage training strategy is adopted. The results show that PINNs can provide satisfactory velocity results and can act as a physics-based denoiser to smooth out the noise on the measured data.
Figure 16: Applying the HFM framework to reconstruct puffing pool fires. Reprinted from Reference 306.
@@
Liu et al. 307 applied PINNs to solve forward problems of 1D and 2D premixed combustion with temperature-dependent properties and a one-step global mechanism of CH4. To the best of our knowledge, although some simplifications are adopted for the transport models and the reaction mechanism is simple, this is the first study on PINNs for combustion reacting flows with detailed physical models and chemical kinetics based on the laws of mass action and Arrhenius. More importantly, after validating the model accuracy on 1D and 2D problems, a surrogate model was constructed for the 2D combustion, where five parameters were added to the NN inputs so that the model can learn multi-condition solutions, as shown in Figure 17. The parameterized PINN surrogate model was shown to accurately predict the fields, be applicable to sensitive analysis, and be much more efficient for industrial designs than traditional numerical simulations. Song et al. 308 applied PINNs to 2D laminar premixed flames considering detailed chemistry of CH4, which was accomplished by integrating the flamelet/progress variable (FPV) model. The equations of mass, momentum, $Z$, and $c$ were solved by PINN, during which the thermophysical properties and source term of $c$ were extracted from the pre-tabulated FPV library, as illustrated in Figure 18. The model can accurately solve the fields with and without observation data, demonstrating the feasibility of PINNs to integrate detailed reaction mechanisms using tabulation methods. Using pseudo-time stepping PINNs (PTS-PINNs) 215, Cao et al. 309 solved the premixed and non-premixed flames with a one-step mechanism of CH4 and obtained satisfactory results. Vanilla PINNs failed to solve the premixed combustion with large gradients and tended to predict non-reacting solutions, but PTS-PINNs showed great convergence performance not only for the premixed combustion but also for non-premixed combustion where the thin reaction zones introduced a higher level of complexity.
Figure 17: Schematic of the parameterized combustion field surrogate model based on PINNs. Reprinted from Reference 307, with permission.
Figure 18: Schematic of the PINNs coupled with the FPV model. Reprinted from Reference 308, with permission.
@@
The above studies only involve forward problems or inverse problems of field reconstruction without parameter inference, which is another important aspect of inverse problems. In computational 1D laminar flames, which are important for the study of flame limit, stretch, laminar flame speed ($s_{L}$), and flamelet tabulation, unknown eigenvalues exist even in forward problems, such as $s_{L}$ in freely-propagating premixed flames 4, 7, 8. Wu et al. 115 established the FlamePINN-1D framework to solve both forward and inverse problems of 1D laminar flames in a unified manner, as illustrated in Figure 19. Three cases were solved: freely-propagating premixed flames with simplified and detailed physical models, and counterflow premixed flames with detailed physical models. The detailed physical models are the same as those in Section (3.1) to (3.5) with a one-step global mechanism of CH4 being considered, except that the governing equations are for 1D reacting stagnation flows. A warmup pretraining strategy and some hard physical constraints were proposed. For example, the predicted species mass fractions are transformed by the Softmax function:
@@
where $Y_{k}^{\#}$ is the output of NN. The transformation ensures that the range constraint ($0\leqslant Y_{k}\leqslant 1$) and species conservation ($\sum_{k}Y_{k}=1$, Equation (3.1)) can be exactly satisfied, so it can be used for various combustion problems. $Y^{\#}$ can be viewed as the logarithmic normalization of $Y$ 271, which can mitigate the multiscale of their magnitudes. The results show that for forward problems, FlamePINN-1D can accurately solve the flame fields and simultaneously infer the $s_{L}$ of freely-propagating premixed flames. For inverse problems, FlamePINN-1D can accurately reconstruct the continuous fields and infer the transport properties and chemical kinetics parameters from noisy sparse observations, which indicates the possibility of optimizing chemical mechanisms from observations of laboratory 1D flames. In fact, in their inverse problems where the activation energy was unknown, the chemical kinetics part can be considered as a CRNN with other parameters are known. It is proven that the model tends to get unburned solutions without warmup pretraining, the similar situation to Reference 309. Other adopted strategies, such as hard constraints, thin-layer normalization 88, and modified MLP with random weight factorization (MMLP-RWF) 82, have also been proven to be essential for the model performance.
Figure 19: FlamePINN-1D for solving forward and inverse problems of 1D laminar flames. (a) Schematic of the FlamePINN-1D framework. (b) Case 1: freely-propagating premixed flames with simplified physical models. (c) Case 2: freely-propagating premixed flames with detailed physical models. (d) Case 3: counterflow premixed flames with detailed physical models. Reprinted from Reference 115, with permission.

### v3

@@ 4.3 Neural ODEs: surrogate models and kinetics discovery
Compared to discrete approaches, NODEs can better learn the continuous dynamics of a system, thus showing better reconstruction and extrapolation performance 291. Since NODE is proposed for learning the first-order dynamics of sequential systems, it is natural to apply it to chemical kinetic systems. Extensive studies have applied NODEs to the modeling of chemical kinetics for both building surrogate models and discovering kinetics (some are shown in Figure 7), where many of them have informed/embedded physics to the NODEs.
Figure 7: Neural ODEs for continuous surrogate modeling and discovering kinetics. (a) The ChemNODE approach 57. $\bm{\psi}$ indicates the thermochemical scalars. (b) The structure of ChemKAN 283. (c) GRNN architecture with embedded stoichiometry and thermodynamics 122. (d) CRNN illustrated for a reaction system with four species and four reaction steps 78. Figures are reprinted with permission.
@@ 4.3.1 Continuous surrogate models
As mentioned in Section 4.2, NODEs can be viewed as continuous surrogate models of chemical kinetics following the paradigm of Equation (4.14). Owoyele et al. 57 first applied NODEs to model chemical kinetics, termed as ChemNODE and shown in Figure 7(a). To improve the training stability, a progressive training strategy was employed, where the species were trained sequentially. For a homogeneous hydrogen-air reactor, ChemNODE showed accurate results across various operating conditions and yielded a speedup of 2.3x over the full chemical mechanism method. Kim et al. 292 used NODEs to learn stiff ODEs. The tests on the ROBER and POLLU problems demonstrated that the scaling of equations and loss functions was essential for NODEs to learn stiff ODEs.
@@
Although the vanilla NODEs are purely data-driven, their interpretability can be enhanced by using the KANs as the network architecture. Koenig et al. 294 proposed the KAN-ODEs, combining the advantage of NODEs that do not require prior knowledge and the advantages of KANs that have increased interpretability and faster neural scaling laws compared to MLPs. The KAN-ODE was validated on learning Lotka-Volterra ODEs, modeling source terms of the Fisher-KPP PDE, and learning 1D PDEs. The scaling performance was shown superior to MLPs and the model performance could be further enhanced by sparsification and using symbolic regression. Subsequently, based on the KAN-ODEs and LeanKAN 295, Koenig et al. 283 proposed the ChemKAN for modeling and acceleration of combustion chemical kinetics. The physics was informed in two ways. First, the architecture strongly mimicked the chemical kinetics, as illustrated in Figure 7(b). A KAN was first used to represent the species production rates (core structure), then they were linearly combined to represent the temperature rate (thermodynamic superstructure) with another KAN to reflect the thermophysical parameter variation. This serial architecture offers the flexibility that one can only learn the species when temperature is not a variable. Second, an optional loss function for element conservation was introduced, which is similar to the PC-NODE and Phy-ChemNODE. The first test case was a biodiesel production modeling, where the thermodynamic superstructure and physical loss were "turned off". Compared with DeepONet, ChemKAN exhibited stronger scaling behaviour on test sets, less tendency toward overfitting, and stronger robustness to noise. In the second case of H2-air combustion, the ability of ChemKAN for computation acceleration was validated. Compared to MLP-based ChemNODE, ChemKAN showed similar speed-up ratios, but solved more species.
@@
In addition to inform physics in the form of loss functions, some studies embed the information of stoichiometry and/or rate expressions into the NODEs to enhance the physical consistency. According to Equation (3.29), the rates of species and reactions are linearly related by the stoichiometric matrix. Thus, if the stoichiometric matrix is known, one can use NODEs to directly predict $\mathbf{RoP}$ and then transform it to $\dot{\bm{\upomega}}$, or vice versa. Based on the known information, Sorourifar et al. 120 developed two versions of physics-enhanced NODEs for chemical reaction systems: r-NODE and k-NODE. The r-NODE approximates $\mathbf{RoP}$ when only the stoichiometry is known and embedded, while the k-NODE approximates the kinetic constants when the mass-action parts of $\mathbf{RoP}$ are further known and embedded. Results suggest that a ground truth model that accurately reflects the data can be learned only when the embedded information accurately reflects the underlying mechanics. Fedorov et al. 121 developed the kinetics-constrained NODEs (KC-NODEs) for kinetic modeling, where the kinetic and thermodynamic knowledge were embedded into the NN architecture, including the stoichiometric matrix, the law of mass action, and the Arrhenius dependency of the rates on temperature. The KC-NODE was successfully applied to the kinetic modeling of the hydrogenation of CO2 to methane using numerical data and to higher hydrocarbons using real experimental data. Kircher et al. 122, 123 proposed the global reaction neural network (GRNN) to learn kinetics from reactor data, where the stoichiometry and thermodynamics were explicitly embedded as the network layers, as shown in Figure 7(c). The GRNN can be trained directly when source data are provided and trained in a NODE way when only reactor data are available, which is a 1D plug flow reactor model in the study. It was demonstrated that the embedded knowledge enhanced the model robustness to significant noise in the data.
@@ 4.3.2 Kinetics discovery
In terms of embedding the knowledge of chemical kinetics into NNs, the chemical reaction neural network (CRNN), proposed by Ji et al. 78 for kinetics discovery, is an earlier work than References 120, 121, 122 and is also trained in a NODE-way. As shown in Figure 7(d), a neuron of CRNN represents a single-step reaction, where the law of mass action (Equation (3.28)) and Arrhenius law (Equation (3.32)) are embedded as the nodes and weights. For the input layer, the weights represent the reaction orders and the bias represents the rate constant in the logarithmic scale. The weights of the output layer correspond to the stoichiometric coefficients. A multi-step reaction system can be represented by stacking the neurons to formulate a CRNN with one hidden layer, where the number of hidden neurons is equal to the number of reactions. Due to the physical interpretability, once the CRNN has been trained, the kinetic parameters can be easily recovered by checking the weights and biases, then the reaction pathways can be fully revealed. The number of hidden nodes/reactions, which is a hyperparameter, can be determined by grid searching and a threshold pruning strategy is suggested to clip the NN weights below a certain threshold 78. The CRNN was first tested on three cases with increasing complexity: an elementary reaction network without temperature dependence, biodiesel production with temperature dependence, and an enzyme reaction network in which some species are present in both reactants and products. Results demonstrated that CRNN could accurately learn the reaction pathways and kinetic parameters from noisy data and even from incomplete data with missing species.
@@ 5 PIML for combustion reacting flows
Choi 302, 2022.1 3D+t Inverse Mass, $\mathbf{U}$, $T$, $Y_{1-3}$ No Van de Vusse Sim. MLP
@@
Sitte 303, 2022.8 2D+t Inverse Mass, $\mathbf{U}$ No Absent Sim. MLP
Liu 304, 2023.9 1D, 2D, 2D+5 Forward Mass, $\mathbf{U}$, $T$, $Y_{1-5}$ No 1-step CH4 Sim. MLP
Song 305, 2024.10 2D Both Mass, $\mathbf{U}$, $Z$, $c$ No Detailed CH4 (FPV table) Sim. MLP
Cao 306, 2024.11 2D Forward Mass, $\mathbf{U}$, $T$, $Y_{1-5}$ No 1-step CH4 Sim. MLP
@@
Liu 307, 2023.12 3D+t Inverse Mass, $\mathbf{U}$, $T$, $_{1-6}, Y\sum Y$ Direct 2-step CH4 Sim. MLP
Liu 308, 2025.9 3D+t Inverse $\mathbf{U}$ Direct Absent Sim. MLP
@@
Peng 309, 2024.4 2D+t Inverse Mass, $\widetilde{\mathbf{U}}$, $\widetilde{T}$, $k$, $\varepsilon$ Mean Absent Sim. MLP
Cao 310, 2025.9 2D, 2D+1 Both Mass, $\widetilde{\mathbf{U}}$, $\widetilde{T}$, $\widetilde{Y}_{1-5}$, $k$, $\varepsilon$, $\mu_{t}$ Mean EDM Sim. MLP
Taassob 311, 2023.10 2D+1 Inverse Mass, $\widetilde{\mathbf{U}}$, $\widetilde{\mathrm{PC}}_{1-2}$, $\widetilde{Z}$, $\mu_{t}$ Mean (learned) Learned Exp. MLP
@@
Yadav 312, 2024.12 (1) 1D (2)(3) 2D Inverse Mass, (1) $c$ ($\mathbf{U}$) (2) $c$ (3) $\widetilde{c}$ (1)(2) No (3) Mean (learned) (1)(2) Pre-learned (3) EBU model (1)(2) Sim. (3) Exp. MLP
Wang 313, 2023.7 1D+t Inverse Mass, $u$, $E$, $c$ No Simplified FRC Sim. MLP
Deng 314, 2024.7 2D+t+2 Inverse Mass, $\widetilde{\mathbf{U}}$ Mean Absent Sim. ResUNet
Zhang 315, 2025.9 2D, 2D+t Inverse Mass, $\mathbf{U}$, $T$, $_{1-5}, Y\sum Y$ (1) No (2) Mean (learned) 1-step CH4 and learned (1) Sim. (2) Exp. MRF-CNN
@@ 5.2 Laminar combustion
This subsection will present a chronological overview of studies on PINNs for laminar subsonic combustion, some of which are illustrated in Figure 8. Choi et al. 302 used PINNs to reconstruct the 3D transient fields of a continuous stirred tank reactor (CSTR) with van de Vusse reaction. The governing equations of two reference frames were considered to enable the reconstruction of the multi-reference frame system. The fields can be accurately reconstructed from the data of velocity, temperature, and species, and some proposed strategies, such as the similarity-based sampling, have been shown to improve the model performance. Wang et al. 263 used PINNs to forward solve the $Z$-equation for a jet diffusion flame and obtained accurate results. Because the assumptions of unity Lewis number and infinite fast chemistry are adopted, the $Z$-equation is free of reaction source terms and the temperature and species can be determined solely by $Z$. Sitte et al. 303 used PINNs to reconstruct the velocity fields in puffing pool fires based on the observations of density, pressure, and temperature, following the HFM framework 247, as illustrated in Figure 8(a). The continuity and momentum equations are taken as the physical losses and a two-stage training strategy is adopted. The results show that PINNs can provide satisfactory velocity results and can act as a physics-based denoiser to smooth out the noise on the measured data.
Figure 8: PINNs for laminar combustion. (a) Applying the HFM framework to reconstruct puffing pool fires 303. (b) Parameterized combustion field surrogate model based on PINNs 304. (c) PINNs coupled with the FPV model 305. (d) FlamePINN-1D for solving forward and inverse problems of 1D laminar flames 115. Figures are reprinted with permission.
@@
Liu et al. 304 applied PINNs to solve forward problems of 1D and 2D premixed combustion with temperature-dependent properties and a one-step global mechanism of CH4. To the best of our knowledge, although some simplifications are adopted for the transport models and the reaction mechanism is simple, this is the first study on PINNs for combustion reacting flows with detailed physical models and chemical kinetics based on the laws of mass action and Arrhenius. More importantly, after validating the model accuracy on 1D and 2D problems, a surrogate model was constructed for the 2D combustion, where five parameters were added to the NN inputs so that the model can learn multi-condition solutions, as shown in Figure 8(b). The parameterized PINN surrogate model was shown to accurately predict the fields, be applicable to sensitive analysis, and be much more efficient for industrial designs than traditional numerical simulations. Song et al. 305 applied PINNs to 2D laminar premixed flames considering detailed chemistry of CH4, which was accomplished by integrating the flamelet/progress variable (FPV) model. The equations of mass, momentum, $Z$, and $c$ were solved by PINN, during which the thermophysical properties and source term of $c$ were extracted from the pre-tabulated FPV library, as illustrated in Figure 8(c). The model can accurately solve the fields with and without observation data, demonstrating the feasibility of PINNs to integrate detailed reaction mechanisms using tabulation methods. Using pseudo-time stepping PINNs (PTS-PINNs) 215, Cao et al. 306 solved the premixed and non-premixed flames with a one-step mechanism of CH4 and obtained satisfactory results. Vanilla PINNs failed to solve the premixed combustion with large gradients and tended to predict non-reacting solutions, but PTS-PINNs showed great convergence performance not only for the premixed combustion but also for non-premixed combustion where the thin reaction zones introduced a higher level of complexity.
@@
The above studies only involve forward problems or inverse problems of field reconstruction without parameter inference, which is another important aspect of inverse problems. In computational 1D laminar flames, which are important for the study of flame limit, stretch, laminar flame speed ($s_{L}$), and flamelet tabulation, unknown eigenvalues exist even in forward problems, such as $s_{L}$ in freely-propagating premixed flames 4, 7, 8. Wu et al. 115 established the FlamePINN-1D framework to solve both forward and inverse problems of 1D laminar flames in a unified manner, as illustrated in Figure 8(d). Three cases were solved: freely-propagating premixed flames with simplified and detailed physical models, and counterflow premixed flames with detailed physical models. The detailed physical models are the same as those in Section (3.1) to (3.5) with a one-step global mechanism of CH4 being considered, except that the governing equations are for 1D reacting stagnation flows. A warmup pretraining strategy and some hard physical constraints were proposed. For example, the predicted species mass fractions are transformed by the Softmax function:
@@
where $Y_{k}^{\#}$ is the output of NN. The transformation ensures that the range constraint ($0\leqslant Y_{k}\leqslant 1$) and species conservation ($\sum_{k}Y_{k}=1$, Equation (3.1)) can be exactly satisfied, so it can be used for various combustion problems. $Y^{\#}$ can be viewed as the logarithmic normalization of $Y$ 271, which can mitigate the multiscale of their magnitudes. The results show that for forward problems, FlamePINN-1D can accurately solve the flame fields and simultaneously infer the $s_{L}$ of freely-propagating premixed flames. For inverse problems, FlamePINN-1D can accurately reconstruct the continuous fields and infer the transport properties and chemical kinetics parameters from noisy sparse observations, which indicates the possibility of optimizing chemical mechanisms from observations of laboratory 1D flames. In fact, in their inverse problems where the activation energy was unknown, the chemical kinetics part can be considered as a CRNN with other parameters are known. It is proven that the model tends to get unburned solutions without warmup pretraining, the similar situation to Reference 306. Other adopted strategies, such as hard constraints, thin-layer normalization 88, and modified MLP with random weight factorization (MMLP-RWF) 82, have also been proven to be essential for the model performance.